OpenStudy (anonymous):

Prove by mathematical induction that 2 + 4 + 6 + 8 + ... + 2n = n^2 + n, for all n >= 1.

OpenStudy (anonymous):

Let the aforementioned be the induction hypothesis. Now consider \(n=1\). Thus, by the induction hypothesis yields the following.\[2=(1)^2+1\]This is obviously true. The induction step is then as follows. Now, let us add \(2(n+1)\) (an operation which is valid for all \(n\ge1\)) to both sides of the induction hypothesis which we will follow with some slight algebraic rearrangement.\[\begin{align} 2+4+6+8+\cdots+2n+2(n+1)&=n^2+n+2(n+1) \\ 2+4+6+8+\cdots+2(n+1)&=(n+1)^2+(n+1) \end{align}\]This demonstrates the induction hypothesis holds for all \(n+1\) when \(n\ge 1\) and thus since the induction hypothesis holds for \(n=1\), the proposed equality holds for all \(n\ge1\).

OpenStudy (anonymous):

Thanks!